define the relation on A={1,2,3} such that aRea of a=b+1 and find it's domain and range


1
Expert's answer
2021-03-10T11:30:56-0500

A= {1,2,3}

R= {(a,b):a=b+1}

Here, a=b+1

"\Rightarrow" b=a-1

So,

When, a=1, b=0

a=2, b=1

a=3, b=2

Now, R= { (1,0), (2,1), (3,2) }

Domain = {1,2,3}

Range = {0,1,2}